A tank contains 60 gallons of 50% solution of glycerin and water. What volume of water must be added to the solution to reduce glycerin concentration to 12%?
Answer by mananth(16949)

--------percent ---------------- quantity
glycerin 50 ---------------- 60 gallons
water 0 ---------------- x gallons
Mixture 10 ---------------- 60 + x gallons

50 * 60 + 0 x = 10 ( 60 + x )
3000 + 0 x = 600 + 10 x
0 x + -10 x = 600 - -3000
-10 x = -2400
/ -10
x = 240
240 gallons of water